OpenStudy (anonymous):

Solve: 6(x - 2) + 3(x - 2) = 18.

OpenStudy (anonymous):

Answers: x = 22/9 x = 2 x = 4

OpenStudy (anonymous):

Apply BIMDAS to get the right answer

OpenStudy (anonymous):

Distribute the 6 to (x-2) so you would get 6x -12 and distribute 3 to (x- 2) so you would get 3x-6 combine like terms and solve.

OpenStudy (anonymous):

First, you want to expand everything out: 6x-12+3x-6=18 Now simplify: 6x+3x-12-6=18 => 9x-18=18 Solve for x: 9x=36 x=4